Two of your classmates have an idea for a new food truck called Funk Seoul Brother, that sells Korean street food and sushi burritos. It would be the only supplier of this food in Athens. Market research suggests that the demand curve for the food will be P=40-(Q/80). It will cost $2000 to set up the food truck, which is necessary before any food can be made. In addition to the setup cost, there is a marginal cost of $6 per entrée for every entrée made.

a. What is the profit-maximizing price and quantity? What profit will this food truck generate?

b. If this market were instead perfectly competitive, what would be the equilibrium price and quantity?

c. How much deadweight loss occurs under the monopolist?
Answer #1

Data Old Powe P = 40 - Q fined cast = 2000 Marginal cost = $6. I = pq - cost 77 = 140-09) 0 - 69-2000 an = 40-8-6=0 dato =) 3

Date/ ROD PONE 4 of market was perfectly P=MC competitive. 6= 40 - 8 80 Q = 34 $_ 34x80 10 = 2720 DWL = (23-6) (2720 -360) DW